Let y vary inversely as x. If x = 2.6hen y = 5, what is y when x = 3.25
Found 2 solutions by stanbon, Nate:
Answer by stanbon(75887)

inversely means y=k/x.
We need to determine "k".
If x=2 then y=5, therefore 5=k/2
So, k=10
EQUATION:
y=10/x
If x=3.25, find y.
y=10/3.25
y= 3.08
Cheers,
Stan H.

Answer by Nate(3500)

Let y vary inversely as x. If x = 2.6 and y = 5, what is y when x = 3.25
-------------------------------------------------------------------------------
if y varies inversely to x....then: yx=yx
xy=xy original
(2.6)(5)=y(3.25)
13=3.25y
4=y